OpenStudy (anonymous):

A store has a special offer on a pack of 12 pencils, 4 pens, and 8 colored markers. The special price is $12.50 for the pack. What fraction of the tools are pencils?

OpenStudy (anonymous):

We know that the total number of items is\[12+4+8=24\] And we know that the number of pencils is 12. So: The fraction of pencils is: \[\frac{12}{24}\] or in simplist terms \[\frac{1}{2}\]

OpenStudy (anonymous):

so if it is 1/2 what is the percent munky99 cuse the teacher may want to know the fraction and percent so write both